OpenStudy (anonymous):

I have a question about chords in a circle I have to draw it for you though. (find x)

OpenStudy (anonymous):



OpenStudy (unklerhaukus):

any arc lengths given?

OpenStudy (anonymous):

No, only description is: "The drawing to the right shows two chords intersecting inside the circle."

OpenStudy (kinggeorge):

Well, we know from the properties of chords that \(5*12 = 60=7.5*x\) so \(x= 8\)

OpenStudy (anonymous):

I did not know that. Will you show me?

OpenStudy (kinggeorge):

I found it under one of the properties here. The third from last property listed under Chord http://en.wikipedia.org/wiki/Circle#Chord

OpenStudy (kinggeorge):

for a nice example that demonstrates a slightly more complicated problem, see http://www.algebralab.org/Word/Word.aspx?file=Geometry_CirclesAnglesChords.xml

OpenStudy (anonymous):

Thanks kinggeorge

OpenStudy (kinggeorge):

your welcome